# The broad-lined Type Ic supernova 2020lao reveals an energetic explosion with no central-engine signatures

### v1

@@
11institutetext: Department of Physics and Astronomy, Aarhus University, Ny Munkegade 120, DK-8000 Aarhus C, Denmark 11email: max@phys.au.dk 22institutetext: National Astronomical Observatory of Japan, National Institutes of Natural Sciences, 2-21-1 Osawa, Mitaka, Tokyo, 181-8588, Japan 33institutetext: Graduate Institute for Advanced Studies, SOKENDAI, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 44institutetext: School of Physics and Astronomy, Monash University, Clayton, VIC 3800, Australia 55institutetext: Astrophysics Research Institute, Liverpool John Moores University, IC2, Liverpool Science Park, 146 Brownlow Hill, Liverpool L3 5RF, UK 66institutetext: The Oskar Klein Centre, Department of Physics, Stockholm University, AlbaNova, 10691 Stockholm, Sweden 77institutetext: Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Drive, Honolulu, HI 96822, USA 88institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), E-08034 Barcelona, Spain 99institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, E-08193 Barcelona, Spain 1010institutetext: Planetary Science Institute, 1700 East Fort Lowell Road, Suite 106, Tucson, AZ 85719-2395, USA 1111institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218-2410, USA 1212institutetext: INAF - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 1313institutetext: Department of Physics, Florida State University, 77 Chieftain Way, Tallahassee, FL, 32306, USA 1414institutetext: INAF - Astrophysics and Space Science Observatory, Via P. Gobetti 101, 40129 Bologna, Italy 1515institutetext: School of Physics and Astronomy, University of Leicester, University Road, Leicester LE1 7RH, UK
The broad-lined Type Ic supernova 2020lao reveals an energetic explosion with no central-engine signatures
@@
(Received date December 3, 2025/ Accepted date )
We present infant-phase observations of the broad-lined Type Ic supernova (SN) 2020lao, including optical spectroscopy beginning within about 48 hours of the inferred explosion epoch and extending to nearly 100 days. The explosion time is constrained by power-law fits to the rising TESS and ZTF light curves, with the first ZTF detection occurring only $\sim 27$ hours after explosion. The optical light curves show a rapid rise of $\approx 8.8$ days and a peak luminosity typical of SNe Ic-BL (i.e., $M_{r}\simeq-18.5$ mag). Unlike some engine-driven SNe Ic-BL events, the early light curve of SN 2020lao shows no evidence of an optical afterglow or excess emission, and the absence of any detectable shock–cooling component in the TESS/ZTF data constrains the progenitor to a compact Wolf-Rayet like star with $R_{\star}\lesssim$ few $R_{\odot}$, ruling out any extended envelope. The spectra resemble those of the X-ray flash associated SN 2006aj but with systematically higher expansion velocities. From Arnett-type fits to the bolometric light curve and measured Fe ii $\lambda$5169 line velocities, we infer a 56Ni mass of $0.23\pm 0.03\penalty 10000\ M_{\odot}$, an ejecta mass ($M_{ej}$) of $3.2\pm 0.8M_{\odot}$, and a kinetic energy of $E_{K}\sim(23.1\pm 12.4)\times 10^{51}$ erg, corresponding to a specific kinetic energy of $E_{K}/M_{ej}\approx(7.2\pm 3.5)\times 10^{51}$ erg $M_{\odot}^{-1}$. Spectral synthesis modeling broadly reproduces the photospheric-phase spectra of SN 2020lao and suggest $E_{K}/M_{ej}\approx 4.9\times 10^{51}$ erg $M_{\odot}^{-1}$. SN 2020lao and SN 2006aj synthesized comparable amounts of 56Ni, yet SN 2020lao exhibits $E_{K}/M_{ej}$ values on the order of 5-10 times larger. Published VLA and Swift/XRT non-detections reveal no afterglow emission, imposing stringent limits on relativistic ejecta and dense circumstellar material. Given that SN 2020lao reaches a specific kinetic energy typical of engine-driven SNe Ic-BL, lack of an early optical excess together with the non-detections in the radio and X-ray bands suggest that if a relativistic jet was launched, the explosion must have been viewed far off axis or the jet was choked before breakout. Otherwise, SN 2020lao represents an extreme non-relativistic SN Ic-BL. This underscores the importance of continued infant-phase, multi-wavelength monitoring of these explosions.
@@ 1 Introduction
Historically, the earliest discoveries of exploding stars depended on the detection of high-energy emissions associated with long-duration gamma-ray bursts and X-ray flashes (Galama1998; Campana2006; Pian2006; Soderberg2006; Ashall19). Follow-up observations of such objects can capture an optical and/or radio afterglow accompanied with the emergence of broad-line Type Ic supernovae (hereafter SNe Ic-BL; Kulkarni1998; Patat2001; Soderberg2006). Some of these events are thought to be powered by a central engine, originally proposed in the collapsar framework (Woosley1993; MacFadyen1999; Woosley2002a) and more recently in magnetar-driven models (Mazzali2013; Metzger2015; Suzuki2018; Piro2019), where the collapse of a massive star leaves behind a compact remnant that drives relativistic jets and injects energy into the ejecta. In recent years, an increasing number of SNe Ic-BL have been discovered through optical transient surveys rather than high-energy emissions (e.g., Anand2024). It remains uncertain whether these events produced high-energy emission that went undetected because of observational limitations, lacked such emission owing to unfavorable viewing angles (e.g., MacFadyen1999; Tanaka2007; Mazzali2005), arose from a distinct mechanism such as a “choked” jet in which the central engine fails to drive a relativistic outflow through the progenitor star (e.g., Irwin2019; MacFadyen2001; RamirezRuiz2002; Lazzati2012; Piran2019), or instead reflect a fundamentally non-relativistic explosion.
@@
The advent of robotic telescopes with high-cadence, all-sky imaging capabilities, combined with machine learning techniques for near-real-time transient identification, has ushered in an era where supernova discovery rates have soared to thousands per year, with an increasing number detected within hours to a day of explosion (hereafter, $t_{exp}$). For example, ALeRCE (Automated Learning for the Rapid Classification of Events) identified ZTF20abbplei (hereafter SN 2020lao) within public stream Zwicky Transient Factory (ZTF; Bellm2019) images taken 2020 May 25.41 UT (i.e., JD 2458994.91; Forster2020). As there were already a handful of previous ZTF non-detections, the source was flagged as a rising SN candidate. A week after discovery, Burke2020 reported the candidate to be a bonafide SN Ic-BL.
@@
SN 2020lao was fortuitously located within the TESS (Transiting Exoplanet Survey Satellite) footprint, with observations covering its position prior to discovery. The full TESS light curve was presented by Vallely2021, who inferred $t_{exp}$ of 2020 May 25.13 UT (i.e., JD $2458994.63\pm 0.06$). More recently, Anand2024 estimated a $t_{exp}$ value of JD $2458993.57\pm 0.75$ based on power-law fits to the early ZTF $g$- and $r$-band photometry. A re-examination of these data sets (see below) yields a $t_{exp}$ value a day earlier, that is on JD $2458993.64^{+0.23}_{-0.20}$, which is adopted throughout this work.
@@
In this study, we analyze optical broadband photometry and spectroscopy of SN 2020lao beginning within 48 hours after $t_{\rm exp}$. These constitute a rare dataset of an infant SN Ic-BL that showed no evidence of high-energy or radio emission (Corsi2023; Anand2024) and no associated optical afterglow component (Burke2020). Despite the lack of such evidence for SN 2020lao, its light curve shape, luminosity, and spectra resemble the X-ray flash–associated SN 2006aj (XRF 060218; Mazzali2006; Pian2006; Sollerman2006), though with spectral lines suggesting an even much more energetic explosion.
@@
Figure 1: An $r$-band image from the Nordic Optical Telescope showing the position of SN 2020lao within its host galaxy, CGCG 169-041.
These similarities, together with the absence of clear engine-driven signatures, raise the question of whether SN 2020lao is related to a GRB viewed off-axis, the outcome of a choked jet, or instead a purely non-relativistic explosion.
@@
Figure 2: The optical BgVriBgVri-band light curves of SN 2020lao, corrected for reddening and vertically offset by arbitrary constants for clarity. The previous ATLAS $o$-band non-detection limit (orange triangle) obtained 0.34-0.34 days prior to $t_{exp}$ (vertical dashed line), is plotted alongside the $r$-band light curve. Additionally, ZTF non-detection limits are included. Segments at the top mark the phases when optical (black) and NIR (red) spectroscopy were obtained (see Table 1).
2 Distance, reddening and host properties
@@
With J2000 coordinates R.A. $17^{h}06^{m}54^{s}.602$ and Decl. $30^{\circ}16^{m}17\aas@@fstack{\prime\prime}35$, SN 2020lao occurred within the face-on, barred spiral galaxy CGCG 169-041 (see Fig. 1). The NASA/IPAC Extragalactic Database (NED) reports a heliocentric redshift of $z=0.031162\pm 0.000117$ (e.g., Marzke1996). Correcting for the Virgo Cluster, Great Attractor, and Shapley supercluster peculiar velocity model (Mould2000) yields a redshift of $z=0.03245$. Using cosmological parameters $\Omega_{m}=0.27$, $\Omega_{\Lambda}=0.73$ (Komatsu2009), and $H_{0}=73.30\pm 1.04$ km s-1 Mpc-1 (Riess2022), this adjusted redshift corresponds to the adopted luminosity distance of $132.7\pm 9.3$ Mpc, or $\mu=35.61\pm 0.15$ mag.
@@
The Milky Way (MW) reddening in the direction of SN 2020lao is $E(B-V)_{MW}=0.044$ mag (Schlafly2011). The host-galaxy reddening appears to be minimal, as indicated by the absence of Na i D in the optical spectra and the post-maximum optical colors (see Sect. 4.2).
@@
We obtained host-galaxy properties through the Blast website (Jones2024_Blast) [^1] https://blast.scimma.org/.. Blast estimates galaxy properties through SED fitting implemented using the Prospector-$\alpha$ model with a non-parametric star formation history (Leja2019; Johnson2021_prospector; Wang2023_prospector, also see Appendix A in Jones2024_Blast for more information). The fitted data include photometry from the Two Micron All Sky Survey (Skrutskie2006), the Wide-field Infrared Survey Explorer (Wright2010), the Sloan Digital Sky Survey (Fukugita1996; York2000), and the Panoramic Survey Telescope and Rapid Response System (Chambers2016) using PSF-matched elliptical apertures.

### v2

# The broad-lined type Ic supernova 2020lao experienced an energetic explosion with no central-engine signatures

@@
11institutetext: Department of Physics and Astronomy, Aarhus University, Ny Munkegade 120, DK-8000 Aarhus C, Denmark 11email: max@phys.au.dk 22institutetext: National Astronomical Observatory of Japan, National Institutes of Natural Sciences, 2-21-1 Osawa, Mitaka, Tokyo, 181-8588, Japan 33institutetext: Graduate Institute for Advanced Studies, SOKENDAI, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan 44institutetext: School of Physics and Astronomy, Monash University, Clayton, VIC 3800, Australia 55institutetext: Astrophysics Research Institute, Liverpool John Moores University, IC2, Liverpool Science Park, 146 Brownlow Hill, Liverpool L3 5RF, UK 66institutetext: The Oskar Klein Centre, Department of Astronomy, Stockholm University, AlbaNova, 10691 Stockholm, Sweden 77institutetext: Institute for Astronomy, University of Hawai’i, 2680 Woodlawn Drive, Honolulu, HI 96822, USA 88institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), E-08034 Barcelona, Spain 99institutetext: Institute of Space Sciences (ICE, CSIC), Campus UAB, Carrer de Can Magrans, s/n, E-08193 Barcelona, Spain 1010institutetext: Planetary Science Institute, 1700 East Fort Lowell Road, Suite 106, Tucson, AZ 85719-2395, USA 1111institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218-2410, USA 1212institutetext: INAF - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy 1313institutetext: Department of Physics, Florida State University, 77 Chieftain Way, Tallahassee, FL, 32306, USA 1414institutetext: INAF - Astrophysics and Space Science Observatory, Via P. Gobetti 101, 40129 Bologna, Italy 1515institutetext: School of Physics and Astronomy, University of Leicester, University Road, Leicester LE1 7RH, UK
The broad-lined type Ic supernova 2020lao experienced an energetic explosion with no central-engine signatures
@@
(Received date / Accepted date )
We present infant-phase observations of the broad-line Type Ic supernova (SN Ic-BL) 2020lao, including optical spectroscopy beginning within about 48 hours of the inferred explosion epoch and extending to nearly 100 days. The explosion time was constrained by power-law fits to the rising TESS and ZTF light curves, with the first ZTF detection occurring only $\sim 27$ hours after explosion. The optical light curves show a rapid rise that lasted for $\approx 8.8$ days and a peak luminosity typical of SNe Ic-BL (i.e., $M_{r}\simeq-18.5$ mag). Unlike some engine-driven SN Ic-BL events, the early light curve of SN 2020lao shows no evidence of an optical afterglow or excess emission, and the absence of any detectable shock–cooling component in the TESS and ZTF data constrains the progenitor to a compact Wolf-Rayet-like star whose $R_{\star}$ is less than or equal to a few times the $R_{\odot}$, ruling out any extended envelope. The spectra resemble those of the X-ray-flash-associated SN 2006aj but with systematically higher expansion velocities. From Arnett-type fits to the bolometric light curve and measured Fe ii $\lambda$5169 line velocities, we infer a 56Ni mass of $0.23\pm 0.03\penalty 10000\ M_{\odot}$, an ejecta mass ($M_{ej}$) of $3.2\pm 0.8M_{\odot}$, and a kinetic energy ($E_{K}$) of $\sim(23.1\pm 12.4)\times 10^{51}$ erg, corresponding to a specific kinetic energy ($E_{K}/M_{ej}$) of $\approx(7.2\pm 3.5)\times 10^{51}$ erg $M_{\odot}^{-1}$. Spectral synthesis modeling broadly reproduces the photospheric-phase spectra of SN 2020lao and suggests $E_{K}/M_{ej}\approx 4.9\times 10^{51}$ erg $M_{\odot}^{-1}$. SN 2020lao and SN 2006aj synthesized comparable amounts of 56Ni, yet SN 2020lao exhibits $E_{K}/M_{ej}$ values on the order of 5-10 times larger. Published VLA and Swift/XRT non-detections reveal no afterglow emission, allowing us to place stringent limits on relativistic ejecta and dense circumstellar material. Given that SN 2020lao reaches a specific kinetic energy typical of engine-driven SNe Ic-BL, the lack of an early optical excess together with the non-detections in the radio and X-ray bands suggests that if a relativistic jet was launched, the explosion must have been viewed far off axis or the jet was choked before breakout. If there was no relativistic jet, SN 2020lao would therefore be an extreme nonrelativistic SN Ic-BL. This underscores the importance of continued infant-phase, multiwavelength monitoring of these explosions.
@@ 1 Introduction
Historically, the earliest discoveries of exploding stars depended on the detection of high-energy emissions associated with long-duration gamma-ray bursts (GRBs) and X-ray flashes (XRFs; Galama et al. 1998; Campana et al. 2006; Pian et al. 2006; Soderberg et al. 2006; Ashall et al. 2019). Follow-up observations of such objects can capture an optical and/or radio afterglow accompanied by the emergence of broad-line Type Ic supernovae (hereafter SNe Ic-BL; Kulkarni et al. 1998; Patat et al. 2001; Soderberg et al. 2006). Some of these events are thought to be powered by a central engine, originally proposed in the collapsar framework (Woosley 1993; MacFadyen and Woosley 1999; Woosley 2002) and more recently in magnetar-driven models (Mazzali et al. 2014; Metzger et al. 2015; Suzuki and Maeda 2018; Piro and Morozova 2019), where the collapse of a massive star leaves behind a compact remnant that drives relativistic jets and injects energy into the ejecta. In recent years, an increasing number of SNe Ic-BL have been discovered through optical transient surveys rather than high-energy emissions (e.g., Anand et al. 2024). It remains uncertain whether these events produced high-energy emission that went undetected because of observational limitations or unfavorable viewing angles (e.g., MacFadyen and Woosley 1999; Tanaka et al. 2007; Mazzali et al. 2005) or arose from a distinct mechanism such as a “choked” jet in which the central engine fails to drive a relativistic outflow through the progenitor star (e.g., Irwin et al. 2019; MacFadyen et al. 2001; Ramirez-Ruiz et al. 2002; Lazzati et al. 2012; Piran et al. 2019), or if they instead reflect a fundamentally nonrelativistic explosion.
@@
The advent of robotic telescopes with high-cadence, all-sky imaging capabilities, combined with machine learning techniques for near-real-time transient identification, has ushered in an era where supernova (SN) discovery rates have soared to thousands per year, with an increasing number detected within hours to a day after the time of the explosion (hereafter, $t_{exp}$). For example, the Automated Learning for the Rapid Classification of Events (ALeRCE) identified ZTF20abbplei (hereafter SN 2020lao) within public stream Zwicky Transient Facility (ZTF; Bellm et al. 2019) images taken 2020 May 25.41 UT (i.e., JD 2458994.91; Forster et al. 2020). As there were already a handful of previous ZTF non-detections, the source was flagged as a rising SN candidate. A week after discovery, Burke et al. (2020) reported the candidate to be a bona fide SN Ic-BL.
@@
SN 2020lao was fortuitously located within the Transiting Exoplanet Survey Satellite (TESS) footprint, with observations covering its position prior to discovery. The full TESS light curve was presented by Vallely et al. (2021), who inferred a $t_{exp}$ of 2020 May 25.13 UT (i.e., JD $2458994.63\pm 0.06$). More recently, Anand et al. (2024) estimated a $t_{exp}$ of JD $2458993.57\pm 0.75$ based on power-law fits to the early ZTF $g$- and $r$-band trytry. A reexamination of these datasets (see below) yields a $t_{exp}$ value a day earlier, that is, on JD $2458993.64^{+0.23}_{-0.20}$, which is adopted throughout this work.
@@
In this study we analyzed optical broadband photometry and spectroscopy of SN 2020lao beginning within 48 hours of $t_{exp}$. This constitutes a rare dataset of an infant SN Ic-BL that showed no evidence of high-energy or radio emission (Corsi et al. 2023; Anand et al. 2024) and no associated optical afterglow component (Burke et al. 2020). To verify these absences, we searched archival high-energy transient catalogs over a broad time window around the inferred explosion epoch and found no reported GRB or X-ray transient spatially and temporally coincident with SN 2020lao. Despite the lack of such high-energy signatures, its light curve shape, luminosity, and spectra closely resemble those of the XRF-associated SN 2006aj (XRF 060218; Mazzali et al. 2006; Pian et al. 2006; Sollerman et al. 2006), albeit with spectral lines suggesting an even more energetic explosion. These similarities, together with the absence of clear engine-driven signatures, raise the question of whether SN 2020lao is related to a GRB viewed off-axis, the outcome of a choked jet, or instead a purely nonrelativistic explosion.
@@
Figure 1: An $r$-band image from the NOT showing the position of SN 2020lao within its host galaxy, CGCG 169-041.
@@
Figure 2: Optical BgVriBgVri-band light curves of SN 2020lao, corrected for reddening and vertically offset by arbitrary constants for clarity. The previous ATLAS $o$-band non-detection limit (orange triangle), obtained 0.34-0.34 days prior to $t_{exp}$ (vertical dashed line), is plotted alongside the $r$-band light curve. ZTF non-detection limits are included. Segments at the top mark the phases when optical (black) and NIR (red) spectroscopy were obtained (see Table LABEL:tab:specjor).
2 Distance, reddening, and host properties
@@
With J2000 coordinates RA $17^{h}06^{m}54^{s}.602$ and Dec $30^{\circ}16^{m}17\aas@@fstack{\prime\prime}35$, SN 2020lao occurred within the face-on, barred spiral galaxy CGCG 169-041 (see Fig. 1). The NASA/IPAC Extragalactic Database (NED) reports a heliocentric redshift of $z=0.031162\pm 0.000117$ (e.g., Marzke et al. 1996). Correcting for the Virgo Cluster, Great Attractor, and Shapley supercluster peculiar velocity model (Mould et al. 2000) yields a redshift of $z=0.03245$. Using cosmological parameters $\Omega_{m}=0.27$, $\Omega_{\Lambda}=0.73$ (Komatsu et al. 2009), and $H_{0}=73.30\pm 1.04$ km s-1 Mpc-1 (Riess et al. 2022), this adjusted redshift corresponds to the adopted luminosity distance of $132.7\pm 9.3$ Mpc, or $\mu=35.61\pm 0.15$ mag.
@@
The Milky Way (MW) reddening in the direction of SN 2020lao is $E(B-V)_{MW}=0.044$ mag (Schlafly and Finkbeiner 2011). The host-galaxy reddening appears to be minimal, as indicated by the absence of Na i D in the optical spectra and the post-maximum optical colors (see Sect. 4.2).
@@
We obtained host-galaxy properties through the Blast website (Jones et al. 2024) [^1] https://blast.scimma.org/.. Blast estimates galaxy properties through spectral energy distribution (SED) fitting implemented using the Prospector-$\alpha$ model with a nonparametric star formation history (Leja et al. 2019; Johnson et al. 2021; Wang et al. 2023; see also Appendix A in Jones et al. 2024 for more information). The fitted data include photometry from the Two Micron All Sky Survey (Skrutskie et al. 2006), the Wide-field Infrared Survey Explorer (Wright et al. 2010), the Sloan Digital Sky Survey (SDSS; Fukugita et al. 1996; York et al. 2000), and the Panoramic Survey Telescope and Rapid Response System (Chambers et al. 2016) using point-spread-function-matched elliptical apertures.